OpenStudy (dls):

A rod of 1000mm and co-efficient of linear expansion alpha=10^-4 /degree is placed symmetrically between fixed walls seperated by 1001mm.The young's modulus of the rod is 10^11 N/m^2 .If the temp is increased by 20 degree/C.then the stress developed in the rod is?

OpenStudy (dls):

\[\LARGE Stress=Y \times strain= Y \times L \alpha \delta T\]

OpenStudy (dls):

So I got 2 * 10^8 which is wrong again \m/

OpenStudy (dls):

@experimentX @shubhamsrg @yrelhan4

OpenStudy (dls):